Find the slope of the line passing through the points (-3, -10) and (4, 5).
Found 2 solutions by fractalier, Boreal:
Answer by fractalier(6550)

Slope m = (y2 - y1) / (x2 - x1)
Here we have
m = (5 - (-10)) / (4 - (-3)) = 15/7

Answer by Boreal(15235)

Change in y/change in x is the slope.
y increases by 15 while x increases by 7.
The slope is 15/7.